# MoEGCL: Mixture of Ego-Graphs Contrastive Representation Learning for Multi-View Clustering

### v3

@@ Introduction
We create the representation of each view by using the Autoencoder architecture (song2018self; zou2023hierarchical; zou2024dai). It contains two parts: an encoder and a decoder. For the $m$-th view, $f^{m}$ represents the encoder. The encoder module generates the representation of each view as follows:
${z}_{i}^{m}=f^{m}\left({x}_{i}^{m}\right),{z}_{i}^{m}\in\mathbb{R}^{d_{\psi}},$ (1)
@@
The $x^{m}_{i}$ is restored by the decoder module utilizing the feature ${z}_{i}^{m}$. Let $g^{m}$ represents the decoder module in the $m$-th view. The recovered sample ${\hat{x}}_{i}^{m}$ is produced by decoding ${z}_{i}^{m}$ in the decoder module:
$\hat{{x}}_{i}^{m}=g^{m}\left({z}_{i}^{m}\right).$ (2)
@@
$\mathcal{L}_{\mathrm{Rec}}=$ $\sum_{m=1}^{M}\left\|{X}^{m}-\hat{{X}}^{m}\right\|_{F}^{2}$ (3)
$=$ $\sum_{m=1}^{M}\sum_{i=1}^{N}\left\|{x}_{i}^{m}-g^{m}\left({z}_{i}^{m}\right)\right\|_{2}^{2},$
@@
First, we create Ego Graph Networks at all views. In particular, we construct the graph $G^{m}=\left\{Z^{m},S^{m}\right\}$ for the $m$-th view, where $Z^{m}=\left[z_{1}^{m};z_{2}^{m};\ldots;z_{N}^{m}\right]$ represents the node features of graph $G^{m}$. $S^{m}$ is the adjacency matrix. In the $m$-th view, the edge between samples $i$ and $j$ is represented by $S_{ij}^{m}$. If $S_{ij}^{m}=1$, samples $i$ and $j$ are connected; otherwise, the two samples are unconnected. We create the adjacency matrix $S^{m}$ using the KNN graph (peterson:73) as follows:
$S_{ij}^{m}=\begin{cases}1&j\in K_{i}^{m};\\ 0&\text{ otherwise.}\end{cases}$ (4)
@@
$S^{m}=(V^{m}_{1};V^{m}_{2};\ldots;V^{m}_{N})^{T},$ (5)
where $V^{m}_{i}$ represents the connection relationship of the ego graph, which is from the $m$-th view of the $i$-th sample. Specifically, it is implemented using the row vectors of the adjacency matrix, also known as the adjacency vector.
@@
We calculate the gating coefficient for each expert with respected to input $z_{i}$. It is implemented through a multi-layer perceptron neural network $mlp^{1}$, followed by classification using a softmax activation function. The number of output categories by softmax is $M$. The specific calculation is as follows:
@@
$S=(V_{1};V_{2};\ldots;V_{N})^{T},$ (9)
where $S$ denotes the fused adjacency matrix.
@@
$\tilde{Z}=\left(\tilde{D}^{-\frac{1}{2}}\tilde{S}\tilde{D}^{-\frac{1}{2}}\right)\left[\left(\tilde{D}^{-\frac{1}{2}}\tilde{S}\tilde{D}^{-\frac{1}{2}}\right)ZW^{0}\right]W^{1},$ (10)
where $\tilde{S}=I_{N}+S$. $I_{N}$ is the identity matrix. $\tilde{D}$ is a diagonal matrix, $\tilde{D}_{ii}=\sum_{j}\tilde{S}_{ij}$. $W^{0}$ and $W^{1}$ are the training parameters. $Z=\{z_{1};z_{2};\ldots;z_{N}\}$ is an input feature, and $\tilde{Z}=\{\tilde{z}_{1};\tilde{z}_{2};\ldots;\tilde{z}_{N}\}$ is an output feature.
@@
We propose the Ego Graph Contrastive Learning (EGCL) module to improve fine-grained graph representation in deep multi-view clustering tasks. EGCL needs to unify the dimensions of each view feature and the fused feature. The specific calculation is as follows:
@@
where the dimension of $\tilde{z}_{i}$ is decreased via the $mlp^{2}$ operator. $d_{\phi}$ denotes the feature dimension of the ${\hat{h}}_{i}$. Similarly, we minimize dimension on each view feature $z^{m}_{i}$ using the $mlp^{3,m}$ operator,
@@
$\mathcal{L}_{\mathrm{Egc}}=-\frac{1}{2N}\sum_{i=1}^{N}\sum_{m=1}^{M}\log\frac{e^{\operatorname{C}\left({\hat{h}}_{i},{h}_{i}^{m}\right)/\tau}}{\sum_{j=1}^{N}e^{(1-{S}_{ij})\operatorname{C}\left({\hat{h}}_{i},{h}_{j}^{m}\right)/\tau}},$ (14)
where $\tau$ denotes the temperature coefficient. Based on Eq. (9), we obtain ${S}_{ij}$. We incorporate ${S}_{ij}$ into Eq. (14) to improve the similarity of the samples in the cluster. When $S_{ij}$ is low, the two samples are not from the same cluster. At this point, the $(1-{S}_{ij})\operatorname{C}({\hat{h}}_{i},{h}_{j}^{m})$ plays a significant role. When $S_{ij}$ is relatively large, it indicates the two samples belong to the same cluster. At this time, the effect of $(1-{S}_{ij})\operatorname{C}({\hat{h}}_{i},{h}_{j}^{m})$ is small.
@@
We utilize the k-means method as the clustering function (mackay2003information; Yan_2023_CVPR). Specifically, the common embedding $\bf{H}$} has the following factorization:
@@
where ${\bf{U}}\in\mathbb{R}^{N\times k}$ denotes the matrix of cluster indicator. We use ${\bf{V}}\in\mathbb{R}^{k\times d^{\phi}}$ as the matrix of clustering center.

### v4

@@ Introduction
We create the representation of each view by using the Autoencoder architecture (song2018self; zou2023hierarchical; zou2024dai). It contains two parts: an encoder and a decoder. For the $m$-th view, $q^{m}$ represents the encoder. The encoder module generates the representation of each view as follows:
${z}_{i}^{m}=q^{m}\left({x}_{i}^{m}\right),{z}_{i}^{m}\in\mathbb{R}^{d_{\psi}},$ (1)
@@
The $x^{m}_{i}$ is restored by the decoder module utilizing the feature ${z}_{i}^{m}$. Let $s^{m}$ represents the decoder module in the $m$-th view. The recovered sample ${\hat{x}}_{i}^{m}$ is produced by decoding ${z}_{i}^{m}$ in the decoder module:
$\hat{{x}}_{i}^{m}=s^{m}\left({z}_{i}^{m}\right).$ (2)
@@
${\Gamma}_{\mathrm{Rec}}=$ $\sum_{m=1}^{M}\left\|{X}^{m}-\hat{{X}}^{m}\right\|_{F}^{2}$ (3)
$=$ $\sum_{m=1}^{M}\sum_{i=1}^{N}\left\|{x}_{i}^{m}-s^{m}\left({z}_{i}^{m}\right)\right\|_{2}^{2},$
@@
First, we create Ego Graph Networks at all views. In particular, we construct the graph $G^{m}=\left\{Z^{m},\Delta^{m}\right\}$ for the $m$-th view, where $Z^{m}=\left[z_{1}^{m};z_{2}^{m};\ldots;z_{N}^{m}\right]$ represents the node features of graph $G^{m}$. $\Delta^{m}$ is the adjacency matrix. In the $m$-th view, the edge between $i$-th sample and $j$-th sample is represented by $\Delta_{ij}^{m}$. If $\Delta_{ij}^{m}=1$, $i$-th sample and $j$-th sample are connected; otherwise, the two samples are unconnected. We create the adjacency matrix $\Delta^{m}$ using the KNN graph (peterson:73) as follows:
$\Delta_{ij}^{m}=\begin{cases}1&j\in K_{i}^{m};\\ 0&\text{ otherwise.}\end{cases}$ (4)
@@
$\Delta^{m}=\begin{pmatrix}V^{m}_{1}\\ V^{m}_{2}\\ \vdots\\ V^{m}_{N}\end{pmatrix}$. (5)
The $i$-th row of $\Delta^{m}$, $V_{i}^{m}$, represents the connection relationship of the ego graph, which is from the $m$-th view of the $i$-th sample. Specifically, it is implemented using the row vectors of the adjacency matrix, also known as the adjacency vector.
@@
We calculate the gating coefficient for each expert with respect to input $z_{i}$. It is implemented through a multi-layer perceptron neural network $mlp^{1}$, followed by classification using a softmax activation function. The number of output categories by softmax is $M$. The specific calculation is as follows:
@@
$\Delta=\begin{pmatrix}V_{1}\\ V_{2}\\ \vdots\\ V_{N}\end{pmatrix}$, (9)
where $\Delta$ denotes the fused adjacency matrix.
@@
$\tilde{Z}=\left(\tilde{D}^{-\frac{1}{2}}\tilde{\Delta}\tilde{D}^{-\frac{1}{2}}\right)\left[\left(\tilde{D}^{-\frac{1}{2}}\tilde{\Delta}\tilde{D}^{-\frac{1}{2}}\right)ZW^{0}\right]W^{1},$ (10)
where $\tilde{\Delta}=I_{N}+\Delta$. $I_{N}$ is the identity matrix. $\tilde{D}$ is a diagonal matrix, $\tilde{D}_{ii}=\sum_{j}\tilde{\Delta}_{ij}$. $W^{0}$ and $W^{1}$ are the training parameters. $Z=\{z_{1};z_{2};\ldots;z_{N}\}$ is an input feature, and $\tilde{Z}=\{\tilde{z}_{1};\tilde{z}_{2};\ldots;\tilde{z}_{N}\}$ is an output feature.
@@
We propose the Ego Graph Contrastive Learning (EGCL) module to improve fine-grained graph representation in deep MVC tasks. EGCL needs to unify the dimensions of each view feature and the fused feature. The specific calculation is as follows:
@@
where the dimension of $\tilde{z}_{i}$ is decreased via the $mlp^{2}$ operator. $d_{\phi}$ denotes the feature dimension of the ${\hat{h}}_{i}$. Similarly, we reduce the dimensionality of each view feature $z^{m}_{i}$ using the $mlp^{3,m}$ operator,
@@
${\Gamma}_{\mathrm{Egc}}=-\frac{1}{2N}\sum_{i=1}^{N}\sum_{m=1}^{M}\log\frac{e^{\operatorname{\xi}\left({\hat{h}}_{i},{h}_{i}^{m}\right)/\tau}}{\sum_{j=1}^{N}e^{(1-{\Delta}_{ij})\operatorname{\xi}\left({\hat{h}}_{i},{h}_{j}^{m}\right)/\tau}},$ (14)
where $\tau$ denotes the temperature coefficient. Based on Eq. (9), we obtain ${\Delta}_{ij}$. We incorporate ${\Delta}_{ij}$ into Eq. (14) to improve the similarity of the samples in the cluster. When $\Delta_{ij}$ is low, the two samples are not from the same cluster. At this point, the $(1-{\Delta}_{ij})\operatorname{\xi}({\hat{h}}_{i},{h}_{j}^{m})$ plays a significant role. When $\Delta_{ij}$ is relatively large, it indicates the two samples belong to the same cluster. At this time, the effect of $(1-{\Delta}_{ij})\operatorname{\xi}({\hat{h}}_{i},{h}_{j}^{m})$ is small.
@@
We utilize the k-means method as the clustering function (mackay2003information; Yan_2023_CVPR). Specifically, the common embedding $\bf{\Phi}$} has the following factorization:
@@
where ${\bf{\Omega}}\in\mathbb{R}^{N\times k}$ denotes the matrix of cluster indicator. We use ${\bf{\Pi}}\in\mathbb{R}^{k\times d^{\phi}}$ as the matrix of clustering center.